# Debiased Sample Selection for Combating Noisy Labels

### v1

@@ B.3 Results on Fine-Grained Noisy Settings
The literature [5] pointed out that noisy label’ generation is strongly associated with fine-grained datasets. Following settings in SNSCL [5], we conduct experiments on four noisy fine-grained datasets, including Stanford Dogs [65] and CUB-200-2011 [66]. A ResNet-18 pre-trained on ImageNet with three experts ($m=3$) is utilized. We synthetically construct two noise types, symmetric and asymmetric noise labels.
@@
The results are shown in Table 7. Our method achieves significant generalization performance even when tackling fine-grained noisy classification and consistently outperforms other methods by a large margin. We consider the MixUp operation to be the primary contribution to the advanced performance on fine-grained datasets since it encourages the model to learn from structured data instead of the unstructured noise [67].
C Experimental Settings
We reported the statistical data of four datasets and the whole experimental setup in Table 8. Besides, the data augmentation strategy for different datasets can be divided into two parts.
  • For CIFAR-10, CIFAR-100, and CIFAR-10N, randomly cropping and horizontally flipping are adopted.
  • For Clothing-1M, we first resize the image to 256x256 and then randomly crop it to 224x224. Besides, the horizontally flipping is adopted.
Meanwhile, we detailedly exhibit the noise generation process by the noise transition matrix. As shown in Figure 10, given a noise ratio $\rho$, the noise transition matrix $T$ is represented as $T_{ij}(\bm{x})=P({\tilde{y}}=j|y=i)$, where $T_{ij}$ denotes that the label of the instance transits from a clean label $i$ to a noisy label $j$. The generation process of instance-dependent label noise can be found in the work [23, 11].
Figure 10: Visualization of the transition matrix for symmetric (left) and asymmetric (right) noise on a ten classification task when the noise ratio is $40\%$.
C.1 Settings for Learning with Noisy Labels on Imbalanced Datasets
We obey the same manner of imbalanced noise construction from RCAL [62] and conduct experiments on synthetic CIFAR-10 and CIFAR-100. Concretely, we set the batch size as 64 and run 200 epochs for both CIFAR-10 and CIFAR-100. The learning rate is set as 0.02 and reduced by a factor of 10 at {100, 150}-th epoch. A ResNet-32 network with 4 experts is utilized. Besides, for a fair comparison, we adopt the same augmentation policy SimAug [68] as RCAL.

### v3

@@ D.4 Results on Fine-Grained Noisy Labels
Recently, one literature [40] pointed out that noisy label generation is strongly associated with fine-grained datasets. Considering the property of wide existence and strong meaningfulness of noisy fine-grained classification, following settings in SNSCL [40], we conduct experiments on four noisy fine-grained datasets, including Stanford Dogs [15] and CUB-200-2011 [42]. Given the selection criterion GMM-based small loss selection, we fine-tuned a (pre-trained) ResNet-18 with three experts. We synthetically construct two noise types, symmetric and asymmetric noise labels.
@@
The results are shown in Table 9. Our method achieves significant generalization performance even when tackling fine-grained noisy classification and consistently outperforms other methods by a large margin. For example, under Stanford Dogs with 30% asymmetric noise, ITEM improves the best method (label smooth) by more than 4%. We consider the MixUp operation to be the primary contribution to the advanced performance on fine-grained datasets for two reasons, 1) it encourages the model to learn from structured data instead of the unstructured noise [1], and 2) it encourages the model to learn a more generalized decision boundary, which largely benefits the fine-grained classification task.